# Onsager vortex clusters on a sphere

### v1

@@ IV.2 Onset of clustering
Hereafter we choose the $\zeta$-axis in the direction of total angular momentum and hence $\mathbf{L}=L_{\zeta}$.
Taking into account the conversation of the angular momentum, Eq.(12) becomes
@@
where $\alpha$ is the Lagrange multiplier associated with the angular momentum and $\omega\equiv\alpha/\beta$ has the meaning of rotation frequency. Accordingly, the vortex number density turns to be
$n_{\pm}=\exp\left[\mp\beta\psi(\textbf{r})\mp\alpha\cos\theta+\gamma_{\pm}\right].$ (28)
Non-uniform solutions at negative temperatures to Eq. (28) describe Onsager vortex states.
Let us consider a solution to Eq. (28) for the given values of $E$ and $L_{\zeta}$, and a nearby solution $n_{\pm}+\delta n_{\pm}$ at $E+\delta E$ and $L_{\zeta}+\delta L_{\zeta}$. The corresponding changes in the constraints are
00 0 $=\int{\rm d}\Omega\ \delta n_{\pm},$ (29)
@@
Combining the variations of Eqs. (11) and (28) in leading order, we obtain
$\nabla^{2}\delta\psi$ $=-\delta{\sigma}(\textbf{r})$
$=n\psi\delta\beta+n\beta\delta\psi-n_{+}\delta\gamma_{+}+n_{-}\delta\gamma_{-}+n\cos\theta\delta\alpha,$
where $n=n_{-}+n_{+}$ is the total density of vortices, and
$\nabla^{2}=\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta\frac{\partial}{\partial\theta}\right)+\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial\phi^{2}}$ (33)
is the Laplace-Bertrami operator in spherical coordinates.
We consider vortex clustered states near the homogeneous state $n_{\pm}=n_{0}=1/4\pi$. For the homogeneous state $\alpha=0$ and $\sigma=\psi=0$, then Eqs. (29),(30) and (31) become
$0=$ $\delta\gamma_{+}+\delta\gamma_{-},$ (34)
$0=$ $\delta\gamma_{+}-\delta\gamma_{-}-2\beta n_{0}\int{\rm d}\Omega\,\delta\psi,$ (35)
@@
where $\delta\alpha$ is assumed to be the same order as $\delta\psi$. Let us introduce operator $\mathcal{P}$:
$\mathcal{P}\delta\psi$ $\equiv\nabla^{2}\delta\psi-2n_{0}\bigg{[}\beta\delta\psi-n_{0}\beta\int d\Omega\,\delta\psi$ (38)
$-3n_{0}\beta\cos\theta\int d\Omega\cos\theta\delta\psi-\frac{3}{2}\cos\theta\delta L_{\zeta}\bigg{]}.$
Then Eq. (IV.2) becomes a zero mode equation of the operator $\mathcal{P}$, i.e.,
@@
$\beta=-\frac{\ell(\ell+1)}{2n_{0}},$ (42)
and
$b=-\frac{n_{0}\beta\omega}{1+n_{0}\beta}$ (43)
for $\ell\neq 1$. While for $\ell=1$, $b$ can be arbitrarily and we choose that $b=0$ without losing generality. Here ${\rm Y}_{\ell m}$ is the spherical harmonic function. For weekly clustered states, the vorticity field $\sigma=\sigma_{\ell m}\equiv\delta\sigma=\epsilon\ell(\ell+1)\psi_{\ell m}$.
@@
$L_{0}$ $=-2n_{0}\beta\int{\rm d}\Omega\cos\theta\ \psi_{\ell m}-\frac{2}{3}\beta\omega.$ (45)
For $\ell=m=0$, $E_{0}=0$ and $L_{0}=0$, hence only modes with $\ell>0$ are relevant. In the following we focus on rotation free states, namely $\omega=0$, and in this case $\psi_{\ell m}=c_{\ell m}\rm{Re}({\rm Y}_{\ell m})$.
IV.2.1 Sandwich state: clusters with zero angular momentum
In this subsection, we consider clustered states with zero angular momentum $L_{0}=0$. Clearly modes with $\ell=2$ have higher statistical weight (larger entropy) than the modes with $\ell>2$.
For $m=0$,
@@ IV.2.1 Sandwich state: clusters with zero angular momentum
where $c_{20}=\sqrt{E_{0}/3}$. This mode describes the clustering of positive (negative) vortices around the poles and negative (positive) vortices around the equator [Fig.2 (a)]. We refer to this state as sandwich state. The onset of the sandwich state occurs at $\beta=\beta_{c}=-12\pi$.
For $m=1,2$,
$\psi_{22}(\theta,\phi)$ $=c_{22}{\rm Re}({\rm Y}_{22})=\frac{c_{22}}{4}\sqrt{\frac{15}{2\pi}}\cos(2\phi)\sin^{2}\theta,$ (47)
$\psi_{21}(\theta,\phi)$ $=c_{21}{\rm Re}({\rm Y}_{21})=-\frac{c_{21}}{4}\sqrt{\frac{15}{2\pi}}\cos\phi\sin 2\theta,$ (48)
where $c_{22}=c_{21}=\sqrt{2E_{0}/3}$. It happens that these two modes are connected by a $\textrm{SO}(3)$ rotation, i.e., ${\rm Y_{21}}(\mathbf{r}^{\prime})={\rm Y_{22}}(\mathbf{r})$, where $\mathbf{r}^{\prime}=R(-\pi/4,\pi/2,\pi/2)\,\mathbf{r}$ and $R(\theta_{1},\theta_{2},\theta_{3})=R_{\zeta}(\theta_{3})R_{\eta}(\theta_{2})R_{\zeta}(\theta_{1})$ is the $\textrm{SO}(3)$ rotation operator and $\{\theta_{i=1,2,3}\}$ are the Euler angles. Therefore, modes $\psi_{22}$ and $\psi_{21}$ are degenerate and describe the same state. We refer to this state as quadrupole state [Fig.2(b)].
Figure 2: Vorticity $\sigma$ for the sandwich (a), the quadrupole (b) and the dipole (c) state. The extreme values of $\sigma$ are scaled to unity for convenience.
For states with $\ell=2$, the dipole moment defined as
[…]
vanishes. In order to quantify the clustered states with zero angular momentum, it is necessary to introduce the quadrupole tensor
[…]
where $\{\textbf{e}_{i=\xi,\eta,\zeta}\}$ are unit vectors in Cartesian coordinates. The quadrupole moment is defined as
@@
and its eigenvalues are $\epsilon c_{20}\sqrt{36\pi/5}\,\{-1,-1,2\}$. The quadrupole momentum for the sandwich state is
@@
where $\nu=1/2$, $Q_{0}=\sqrt{216\pi/15}$ and $E_{c}=0$. For quadrupole states, the quadrupole tensors are
$\mathcal{Q}[\sigma_{21}]$ $=-\epsilon{c_{21}}\sqrt{\frac{54\pi}{5}}\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{array}\right){,}$ (60)
and
$\mathcal{Q}[\sigma_{22}]$ $=\epsilon{c_{22}}\sqrt{\frac{54\pi}{5}}\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&0\\ \end{array}\right).$ (64)
Since $\sigma_{21}(R\mathbf{r})=\sigma_{22}(\mathbf{r})$, it is easy to see that $\mathcal{Q}[\sigma_{21}]=R\mathcal{Q}[\sigma_{22}]R^{T}$, where $R=R(-\pi/4,\pi/2,\pi/2)$. Their common eigenvalues are $-{c_{21}}\epsilon\sqrt{54\pi/5}\,\{-1,0,1\}$. Moreover the quadrupole momentum for the quadrupole state takes the same value as for the sandwich state, namely, $Q_{q}=Q[\sigma_{21}]=Q[\sigma_{22}]=Q[\sigma_{20}]=Q_{s}$.
Using Eq. (9), we find that the entropy of the sandwich state is larger than the quadrupole state,
$S[\sigma_{20}]>S[\sigma_{21}]=S[\sigma_{22}].$ (65)
Hence under condition of zero angular momentum $L_{0}=0$, the sandwich state has the highest statistical weight. Although this conclusion is made based on the perturbative analysis at low energies, it will be confirmed by later analysis at high energies and Monte Carlo simulations at moderate energies.
IV.2.2 dipole states: vortex clustered states with finite angular momentum
We now consider states with finite angular momentum $L_{0}\neq 0$. For $\ell=1,m=0$, the corresponding stream function is
$\psi_{10}=c_{10}{\rm Y}_{10}=c_{10}\sqrt{\frac{3}{4\pi}}\cos\theta,$ (66)
where $c_{10}=\sqrt{E_{0}}$.
The dipole moment for this state does not vanish and its magnitude reads
$D=|\mathbf{D}|=4\sqrt{\frac{\pi}{3}}\sqrt{E}.$ (67)
We refer this state as dipole state [Fig. 2 (c)] and the onset of the dipole state occurs at $\beta=\beta^{d}_{c}=-4\pi$. For the dipole state, the angular momentum is $L_{\zeta}=4\sqrt{\pi/3}\sqrt{E}$. Note that the vorticity distribution for the dipole state at low energies is identical to the vorticity distribution for a stationary vortex flow on a sphere Xiong and Yu (2024).
IV.3 High energy configurations
@@
In stereographic coordinates Eq. (14) becomes
$\ (1+z\bar{z})^{2}\partial_{z}\partial_{\bar{z}}\psi=e^{\beta\psi(z,\bar{z})+\gamma_{-}}-e^{-\beta\psi(z,\bar{z})+\gamma_{+}}$ (68)
for rotation free states ($\omega=0$). The stereographic projection maps the southern hemisphere onto a disc on the $\zeta=0$ plane (Fig.1) and transforms vortices in the region around southern pole and vortices in the region around the equator to the region around the origin and the region near the boundary of the disc, respectively.
IV.3.1 Sandwich states at high energies
In the high energy limit, the interaction between vortices near the poles and antivortices near the equator can be neglected. Let us firstly consider the region $|z|\ll 1$, in which Eq. (68) can be well approximated as
$\partial_{z}\partial_{\bar{z}}\psi=-e^{-\beta\psi(z,\bar{z})+\gamma_{+}}.$ (69)

### v2

@@ III.2 Continuous limit and the mean-field theory
Hereafter we choose the $\zeta$-axis in the direction of total angular momentum and hence $\bm{L}=L_{\zeta}$ and we set $\rho_{s}=\kappa=R=k_{B}=1$ for convenience. In this convention $\xi\ll\bar{d}\ll 1$.
In the continuous limit, $a_{i}\rightarrow 0$ and $N_{i,\pm}/N_{\pm}a_{i}\rightarrow n_{\pm}(\bm{r})$, and the Eq. (18) gives rise to the thermodynamic entropy per vortex for given vortex distributions $n_{\pm}$ Joyce and Montgomery (1973); Eyink and Sreenivasan (2006):
$S[n_{+},n_{-}]=-\int{\rm d}\Omega\left(n_{+}\log n_{+}+n_{-}\log n_{-}\right),$ (27)
and the most probable vortex density distribution is given by maximizing the entropy Eq. (27) subject to given energy $E$, vortex number $N_{\pm}$ and angular momentum $L_{\zeta}$:
@@
where $\beta$ and $\mu_{\pm}$ are Lagrange multipliers and have the interpretation of inverse temperature and chemical potentials in microcanonical ensemble, respectively. Here $\alpha$ is the Lagrange multiplier associated with the angular momentum and $\omega\equiv\alpha/\beta$ has the meaning of rotation frequency.
The variation equation (28) gives rise to
$n_{\pm}=\exp\left[\mp\beta\psi(\bm{r})\mp\alpha\cos\theta+\gamma_{\pm}\right].$ (29)
[…]
$\nabla^{2}\psi(\bm{r})$ $=$ $\exp[\beta\psi(\bm{r})+\alpha\cos\theta+\gamma_{-}]$
$-\exp[-\beta\psi(\bm{r})-\alpha\cos\theta+\gamma_{+}].$
Non-uniform solutions at negative temperatures to Eq. (III.2) describe Onsager vortex states.
III.3 Onset of clustering
The onset of clustering involves global eigen-modes of the Laplace-Bertrami operator. It is then convenient to use spherical coordinates. Let us consider a solution to Eq. (III.2) for the given values of $E$ and $L_{\zeta}$, and a nearby solution $n_{\pm}+\delta n_{\pm}$ at $E+\delta E$ and $L_{\zeta}+\delta L_{\zeta}$. The corresponding changes in the constraints are
$0$ $=\int{\rm d}\Omega\ \delta n_{\pm},$ (31)
@@ III.3 Onset of clustering
For the homogeneous state $n_{\pm}=n_{0}=1/4\pi$, and $\alpha=\sigma=\psi=0$. Then Eqs. (31),(32) and (33) become
${\rm e}^{-\delta\gamma_{\pm}}$ $=n_{0}\int{\rm d}\Omega\,\exp(\mp\beta\delta\psi+\delta\alpha\cos\theta),$ (34)
@@
$\delta L_{\zeta}$ $=\int{\rm d}\Omega\ \delta\sigma\cos\theta\ .$ (36)
Expanding $\delta\sigma$ and Eq. (31) up to the leading order in $\delta\psi$, we have
$\delta\sigma$ $=-2n_{0}(\beta\delta\psi-\delta\gamma_{+}+\cos\theta\delta\alpha),$ (37)
$\delta\gamma_{+}$ $=-\delta\gamma_{-}=\beta n_{0}\int{\rm d}\Omega\,\delta\psi,$ (38)
and as it follows Eqs. (35) and  (36) become
$\delta E$ $=-n_{0}\int{\rm d}\Omega\,(\beta\delta\psi-\delta\gamma_{+}+\cos\theta\delta\alpha)\delta\psi,$ (39)
@@
where $\delta\alpha$ is assumed to be the same order as $\delta\psi$.
Combining Eqs. (22) and (37), up to leading order in $\delta\psi$, we obtain
$\nabla^{2}\delta\psi$ $=-\delta{\sigma}(\bm{r})=2n_{0}(\beta\delta\psi-\delta\gamma_{+}+\cos\theta\delta\alpha).$ (41)
Let us express the changes of the Lagrange multipliers in terms of the changes of the corresponding conserved quantities and introduce the operator $\mathcal{P}$:
$\mathcal{P}\delta\psi$ $\equiv\nabla^{2}\delta\psi-2n_{0}\bigg[\beta\delta\psi-n_{0}\beta\int{\rm d}\Omega\,\delta\psi$ (42)
$-3n_{0}\beta\cos\theta\int{\rm d}\Omega\cos\theta\delta\psi-\frac{3}{2}\cos\theta\delta L_{\zeta}\bigg].$
Then Eq. (41) becomes a zero mode equation of the operator $\mathcal{P}$, i.e.,
@@
$\beta=\beta_{c,\ell}=-\frac{\ell(\ell+1)}{2n_{0}},$ (46)
and $b=-n_{0}\beta\omega/(1+n_{0}\beta)$ for $\ell\neq 1$. While for $\ell=1$, $b$ can be arbitrary and we choose that $b=0$ without losing generality. Here ${\rm Y}_{\ell m}$ is the spherical harmonic function. For weakly clustered states, the vorticity field $\sigma=\sigma_{\ell m}\equiv\delta\sigma=\epsilon\ell(\ell+1)\psi_{\ell m}$.
@@
$L_{0}$ $=-2n_{0}\beta_{c,\ell}\int{\rm d}\Omega\cos\theta\ \psi_{\ell m}-\frac{2}{3}\beta_{c,\ell}\omega.$ (48)
For $\ell=m=0$, $E_{0}=0$ and $L_{0}=0$, hence only modes with $\ell>0$ are relevant. In the following we focus on rotation free states, namely $\omega=\alpha=\delta\alpha=0$, and in this case $\psi_{\ell m}=c_{\ell m}\rm{Re}({\rm Y}_{\ell m})$.
III.3.1 dipole states: vortex clustered states with finite angular momentum
Let us firstly consider the state with $\ell=1$ and $m=0$, the corresponding stream function is
$\psi_{10}=c_{10}{\rm Y}_{10}=c_{10}\sqrt{\frac{3}{4\pi}}\cos\theta,$ (49)
where $c_{10}=\sqrt{E_{0}}$. This state carries finite angular momentum and $L_{0}\neq 0$.
The dipole moment defined as
$\bm{D}\equiv\int{\rm d}\Omega\ \sigma(\bm{r})\,\bm{r},$ (50)
and for this state the dipole moment reads
$D_{\eta}=D_{\xi}=0\quad\text{and}\quad D_{\zeta}=4\sqrt{\frac{\pi}{3}}\sqrt{E}.$ (51)
We refer to this state as dipole state and the onset of the dipole state occurs at $\beta=\beta_{c,1}\equiv\beta^{d}_{c}=-4\pi$. For the dipole state, the angular momentum $L_{\zeta}=4\sqrt{\pi/3}\sqrt{E}$. It is interesting to note that the vorticity distribution for the dipole state at low energies is identical to the vorticity distribution for a stationary vortex flow on a sphere Xiong and Yu (2024), obtained from the vortex fluid theory.
For $m=\pm 1$, the corresponding stream functions are $\psi_{11}=-\psi_{1,-1}=c_{11}{\rm Re}({\rm Y}_{11})=-c_{11}\sqrt{3/8\pi}\sin\theta\cos\phi$, where $c_{11}=\sqrt{2}c_{10}$. It happens that these modes are connected by a $\textrm{SO}(3)$ rotation, namely, $\psi_{11}(\bm{r})=\psi_{10}(\bm{r}_{1})$ and $\bm{r}_{1}=R(0,\pi/2,0)\,\bm{r}$. Here $R(\theta_{1},\theta_{2},\theta_{3})=R_{\zeta}(\theta_{3})R_{\eta}(\theta_{2})R_{\zeta}(\theta_{1})$ is the $\textrm{SO}(3)$ rotation operator and $\{\theta_{i=1,2,3}\}$ are the Euler angles. Therefore modes $\psi_{10}$, $\psi_{11}$ and $\psi_{1-1}$ describe the same dipole state while with different dipole moment orientations. Hereafter we focus on the dipole state described by the mode $\psi_{10}$ for which the dipole moment is along the $\zeta$-axis [Fig. 2 (a)] .
III.3.2 Sandwich states: clusters with zero angular momentum
In this subsection, we consider clustered states with zero angular momentum $\bm{L}=0$ (or $L_{0}=0$). Hence only modes with $\ell>1$ are relevant and $\delta L_{\zeta}=\delta\alpha=0$. Intuitively modes with larger $\ell$ describe richer and finer structure and should have lower entropy. To ensure it is the case, let us expand $S[n_{0}+\delta n_{-},n_{0}+\delta n_{+}]$ around the uniform density $n_{0}$ up to second order in $\delta n_{\pm}$, and we obtain
[…]
$=$ $S_{0}-\frac{1}{2n_{0}}\left[\left(\int{\rm d}\Omega\ {\rm e}^{-\beta\delta\psi}\right)^{-2}\int{\rm d}\Omega\,{\rm e}^{-2\beta\delta\psi}\right$.
[…]
where the first order term in $\delta n_{\pm}$ vanishes due to the constraint Eq. (31) and $S_{0}=-\int{\rm d}\Omega\ 2n_{0}\log n_{0}$ is the entropy for the homogeneous state. Here we have used $\delta n_{\pm}=n_{0}[{\rm e}^{\delta\gamma_{\pm}}{\rm e}^{\mp\beta\delta\psi}-1]$ and Eq. (34) for $\delta\alpha=0$. To leading order in $\epsilon$, $\delta n_{\pm}\simeq\mp\epsilon n_{0}\beta_{c,\ell}\psi_{\ell,m}$ and $\delta\gamma_{\pm}=0$, then Eq. (52) becomes
$S$ $\simeq S_{0}-\left[n_{0}\beta^{2}_{c,\ell}\int{\rm d}\Omega\,\psi^{2}_{\ell m}\right]\epsilon^{2}$ (53)
$=S_{0}-\frac{1}{E_{0}}\left[n_{0}\beta^{2}_{c,\ell}\int{\rm d}\Omega\,\psi^{2}_{\ell m}\right]\delta E,$
and one can easily show that modes with $\ell=2$ have higher statistical weight (larger entropy) than the modes with $\ell>2$ (see Sec. A in Appendix).
Hence we only need to focus on states with $\ell=2$. For $m=0$,
@@ III.3.2 Sandwich states: clusters with zero angular momentum
where $c_{20}=\sqrt{E_{0}/3}$. This mode describes the clustering of positive (negative) vortices around the poles and negative (positive) vortices around the equator [Fig.2 (b)]. We refer to this state as sandwich state. The onset of the sandwich state occurs at $\beta=\beta_{c,2}\equiv\beta_{c}=-12\pi$. For $m=1$ and $m=2$, $\psi_{22}(\theta,\phi)=c_{22}{\rm Re}({\rm Y}_{22})=(c_{22}/4)\sqrt{15/2\pi}\cos(2\phi)\sin^{2}\theta$, $\psi_{21}(\theta,\phi)=c_{21}{\rm Re}({\rm Y}_{21})=-(c_{21}/4)\sqrt{15/2\pi}\cos\phi\sin 2\theta$, where $c_{22}=c_{21}=\sqrt{2E_{0}/3}$. Here the two modes are also connected by a $\textrm{SO}(3)$ rotation, i.e., $\psi_{21}(\bm{r}_{2})=\psi_{22}(\bm{r})$, where $\bm{r}_{2}=R(-\pi/4,\pi/2,\pi/2)\,\bm{r}$. Therefore, modes $\psi_{22}$ and $\psi_{21}$ describe the same state and we refer to this state as quadrupole state. Note that $\psi_{2,-2}=\psi_{2,2}$ and $\psi_{2,-1}=-\psi_{2,1}$, hereafter we focus on the quadrupole state described by the mode $\psi_{21}$[Fig.2(c)].
Figure 2: Vorticity $\sigma$ for the dipole (a), the sandwich (b) and the quadrupole (c) state. The extreme values of $\sigma$ are scaled to unity for convenience.
For states with $\ell=2$, the dipole moment $\bm{D}$ vanishes. In order to quantify the clustered states with zero angular momentum, it is necessary to introduce the quadrupole tensor
$\mathcal{Q}_{ij}\left[\sigma(\bm{r})\right]\equiv\frac{1}{2}\int{\rm d}\Omega\,\Big[3(\bm{r}\cdot\bm{e}_{i})(\bm{r}\cdot\bm{e}_{j})-\delta_{ij}\Big]{\sigma}(\bm{r}),$ (55)
where $\{\bm{e}_{i=\xi,\eta,\zeta}\}$ are unit vectors in Cartesian coordinates. The quadrupole moment is defined as
@@
and its eigenvalues are $\epsilon c_{20}\sqrt{36\pi/5}\,\{-1,-1,2\}$. The corresponding quadrupole momentum is
@@
where $\nu=1/2$, $Q_{0}=\sqrt{216\pi/15}$ and $E_{c}=0$. For the quadrupole state,
$\mathcal{Q}[\sigma_{21}]$ $=-\epsilon{c_{21}}\sqrt{\frac{54\pi}{5}}\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{array}\right){,}\$(65)
and the eigenvalues of $\mathcal{Q}[\sigma_{21}]$ are ${c_{21}}\epsilon\sqrt{54\pi/5}\,\{-1,0,1\}$. The corresponding quadrupole momentum takes the same value as for the sandwich state, namely, $Q_{q}=Q_{s}$. Note that since $\sigma_{21}(R\bm{r})=\sigma_{22}(\bm{r})$, it is easy to see that $\mathcal{Q}[\sigma_{21}]=R\mathcal{Q}[\sigma_{22}]R^{T}$, where $R=R(-\pi/4,\pi/2,\pi/2)$.
Applying Eq. (53), we find that up to order $\delta E$, the sandwich state and the quadrupole state have the same entropy and the entropy difference appears in order $\delta E^{2}$:
$S[{\rm sandwich}]-S[{\rm quadrupole}]\simeq\frac{5\beta_{c}^{4}E_{0}^{2}\epsilon^{4}}{576\pi^{2}}=\frac{5\beta_{c}^{4}}{576\pi^{2}}(\delta E)^{2}.$
Hence under condition of zero angular momentum, the sandwich state has the highest statistical weight. Although this conclusion is made based on the perturbative analysis at low energies, it will be confirmed by later analysis at high energies and Monte Carlo simulations at moderate energies.
III.4 High energy configurations
@@
In stereographic coordinates Eq. (III.2) becomes
$\ (1+z\bar{z})^{2}\partial_{z}\partial_{\bar{z}}\psi={\rm e}^{\beta\psi(z,\bar{z})+\gamma_{-}}-{\rm e}^{-\beta\psi(z,\bar{z})+\gamma_{+}}$ (67)
for rotation free states ($\omega=\alpha=0$). The stereographic projection maps the southern hemisphere onto a disc on the $\zeta=0$ plane (Fig.1) and transforms vortices in the region around southern pole and vortices in the region around the equator to the region around the origin and the region near the boundary of the disc, respectively.
III.4.1 Sandwich states at high energies
In the high energy limit, the interaction between vortices near the poles and antivortices near the equator can be neglected. For $|z|\ll 1$, Eq. (67) is well approximated as
$\partial_{z}\partial_{\bar{z}}\psi=-{\rm e}^{-\beta\psi(z,\bar{z})+\gamma_{+}},$ (68)
which can be regarded as the mean-field equation of describing distributions of chiral vortices confined to a (flat) disc and is well-defined on the whole disc domain. Hence finding the approximated distribution of vortices near the poles at high energies is boiled down to searching for distributions of chiral vortices confined to a disc with proper boundary conditions.